### v1

# GLIP-OOD: Zero-Shot Graph OOD Detection with Foundation Model

Abstract
Out-of-distribution (OOD) detection is critical for ensuring the safety and reliability of machine learning systems, particularly in dynamic and open-world environments. In the vision and text domains, zero-shot OOD detection—which requires no training on in-distribution (ID) data—has made significant progress through the use of large-scale pretrained models such as vision-language models (VLMs) and large language models (LLMs). However, zero-shot OOD detection in graph-structured data remains largely unexplored, primarily due to the challenges posed by complex relational structures and the absence of powerful, large-scale pretrained models for graphs. In this work, we take the first step toward enabling zero-shot graph OOD detection by leveraging a graph foundation model (GFM). We show that, when provided only with class label names, the GFM can perform OOD detection without any node-level supervision—outperforming existing supervised methods across multiple datasets. To address the more practical setting where OOD label names are unavailable, we introduce GLIP-OOD, a novel framework that employs LLMs to generate semantically informative pseudo-OOD labels from unlabeled data. These labels enable the GFM to capture nuanced semantic boundaries between ID and OOD classes and perform fine-grained OOD detection—without requiring any labeled nodes. Our approach is the first to enable node-level graph OOD detection in a fully zero-shot setting, and achieves state-of-the-art performance on four benchmark text-attributed graph datasets.
**footnotetext: Equal contribution.
1 Introduction
Out-of-distribution (OOD) detection [31, 14, 9] is the task of identifying inputs that deviate significantly from the training distribution. As machine learning models are increasingly deployed in dynamic and safety-critical environments, the ability to reliably detect OOD instances has become important for maintaining robustness, improving trustworthiness, and preventing catastrophic failures. Applications span many domains, including medical diagnosis [23], autonomous driving [6], video analytics [11], and financial anomaly detection [2], where encountering unseen or unexpected data is common and consequential.
Graph OOD Detection and Current Landscape. Recently, OOD detection has gained traction in graph-structured data (known as graph OOD detection [27, 15, 29, 22]), where the goal is to detect anomalous or previously unseen nodes, edges, or subgraphs. It is important to the domains such as social networks [28], citation graphs, and e-commerce platforms, where new entities (e.g., emerging users, novel research topics, or innovative products) may appear during inference. These graph OOD detection methods generally require a substantial number of labeled in-distribution (ID) nodes to train a graph neural network (GNN)-based classifier, which subsequently produces class probability distributions (logits) for all nodes in the graph [27, 15, 29]. OOD detection is then performed in a post-hoc manner by applying scoring functions—such as energy-based [27], maximum softmax probability [9], or uncertainty-based metrics [15]—to these logits. Nodes with high energy or entropy scores are considered likely to be OOD samples. However, these methods are fundamentally limited by their reliance on labeled ID data, which can be expensive and impractical to obtain in real-world graph scenarios. This raises a natural question: Can we perform graph OOD detection without relying on labeled nodes at all?
Figure 1: GLIP-OOD-R denotes using real labels, while GLIP-OOD-L uses LLM-generated labels. When provided with only label names and no node-level supervision, GLIP-OOD achieves OOD detection performance that surpasses all supervised graph learning methods trained on large amounts of ID-labeled data across all datasets.
Zero-Shot OOD Detection in Vision and Text. In computer vision and NLP areas, zero-shot OOD detection—which requires no training on ID data—has emerged as a compelling research direction. In the image domain, methods [24, 18, 5, 1] leverage the capabilities of large-scale vision-language models (VLMs), e.g., CLIP [19], to detect OOD samples across a variety of ID datasets without requiring any labeled training examples. For example, CLIPN [24] enhances CLIP with the ability to distinguish between OOD and ID samples by leveraging both positive semantic prompts and negation-based prompts, enabling zero-shot OOD detection in the image domain. ZOC [5] extends CLIP by dynamically generating candidate unseen labels per test image and introducing a confidence score based on feature-space similarity to both seen and generated unseen labels. In the text domain, large language models (LLMs) have demonstrated impressive zero-shot capabilities by leveraging natural language prompts and contextual understanding to detect OOD content [32, 10]. The ability to perform effectively with limited data makes LLM-based OOD detection particularly attractive for real-world applications. Despite these advances, zero-shot OOD detection in graph domains remains largely unexplored due to the unique challenges posed by relational structures. More importantly, until now, there has been no powerful large-scale foundation model—like VLMs and LLMs—for the graph domain.
Our Proposal: Zero-Shot OOD Detection in Graphs. In this paper, we take the first step toward enabling zero-shot graph OOD detection using a graph foundation model (GFM) [33, 12, 13, 3]. Similar to CLIP, the original GFM lacks built-in OOD detection capabilities—it is designed to match any given node to one of the provided seen class labels. To enable effective OOD detection, we first introduce only the label names to the GFM and directly derive the output logits from it to perform zero-shot OOD detection. As shown in Fig. 1, we empirically find that when provided solely with label names, the GFM achieves OOD detection performance that surpasses all supervised graph learning methods relying on large numbers of ID labels to train an ID classifier across all datasets. This demonstrates that the GFM is inherently capable of performing zero-shot OOD detection and holds significant potential to become highly OOD-aware, provided the ID and OOD classes are clearly defined. However, in practical open-world scenarios, we are often unable to access the actual OOD label space. To address this, we propose generating a set of pseudo-OOD labels based on the inherent knowledge within the graph and performing OOD detection on an augmented label space that includes both the original ID classes and the newly synthesized OOD labels.
In contrast to image and text OOD detection, node-level graph OOD detection typically operates in a transductive setting, where the model has access to the complete set of nodes during training, but only the labels for the ID classes are provided. Within this transductive setting, we propose leveraging LLMs to generate pseudo-OOD label names based on the initially unlabeled graph. Specifically, we begin by randomly sampling a small subset of unlabeled nodes and use the LLM to perform coarse-grained zero-shot OOD detection. The LLM is given the names of the ID classes and asked to determine whether a query node, described by its textual attributes, belongs to any of the known ID categories. If the LLM predicts that the node does not match any ID class, we consider it a pseudo-OOD node. For each such node, we further prompt the LLM to summarize its textual description and generate a representative OOD category name, thereby creating a set of pseudo-OOD labels. We then augment the original ID label set with these pseudo-OOD labels and perform zero-shot OOD detection using the GFM across the expanded label space. This approach enables the GFM to assign each node to either one of the known ID classes or a synthesized OOD category, thereby facilitating fine-grained OOD detection without requiring any real OOD labels or manual annotations.
We summarize our key contributions as follows:
  • To the best of our knowledge, we are the first to study the problem of zero-shot graph OOD detection and to provide a comprehensive setting and framework for enabling node-level OOD detection without any node-level supervision.
  • We propose GLIP-OOD, which uses an LLM to generate pseudo-OOD labels that are semantically distant from the given ID labels but close to the actual unknown OOD labels in the transductive graph setting. We then perform zero-shot OOD detection directly with the GFM in the augmented label space.
  • Experimental results demonstrate that our proposed zero-shot graph OOD detection method outperforms all existing supervised approaches on four benchmark TAG datasets when provided with all label names but without any node-level supervision. Furthermore, when only ID label names are provided, GLIP-OOD achieves OOD detection performance comparable to that of supervised methods, which require training an ID classifier using a large number of labeled ID nodes.
2 Preliminary
Our study focus on OOD detection on TAGs. A TAG is represented as $G_{T}=(\mathcal{V},\mathbf{A},\mathbf{T},\mathbf{X})$. The set of nodes is $\mathcal{V}=\{v_{1},\dots,v_{n}\}$, where each node is associated with raw text attributes $\mathbf{T}=\{t_{1},t_{2},\dots,t_{n}\}$. These text attributes are converted into embeddings $\mathbf{X}=\{x_{1},x_{2},\dots,x_{n}\}$ using a pretrained model [21]. The adjacency matrix $\mathbf{A}\in\{0,1\}^{n\times n}$ encodes graph connectivity, where $\mathbf{A}[i,j]=1$ indicates an edge between nodes $v_{i}$ and $v_{j}$.
Node-level Graph OOD Detection. We have a labeled node set $\mathcal{V}_{I}$ and a large unlabeled node set $\mathcal{V}_{U}$. Each labeled node belongs to one of $K_{\text{ID}}$ known classes $\mathcal{Y}_{\text{ID}}=\{y_{k}\}_{k=1}^{K_{\text{ID}}}$, while an unlabeled node may belong to an unknown class not included in $\mathcal{Y}_{\text{ID}}$. The problem is framed in a semi-supervised, transductive setting, where we can access the full set of nodes during training but only a subset of class labels (ID classes). Our goal is, for each node $v_{i}\in\mathcal{V}_{U}$, determine whether it is from the $K_{\text{ID}}$ ID classes or from the OOD unknown classes.
Zero-shot Graph OOD Detection with All Labels. We provide the model only with the class names and do not assume access to any node labels in the graph. That is, all nodes in the graph are unlabeled, and the model directly outputs an OOD score for each input node based solely on the label names, node attributes $X$, and the adjacency matrix $A$.
Zero-shot Graph OOD Detection. Only the names of ID classes, node attributes $X$, and the adjacency matrix $A$ are available, with no access to any labels for ID or OOD nodes. This setting is consistent with the zero-shot OOD detection paradigm in the image domain.
3 Methodology
We address the zero-shot graph OOD detection problem by adapting GFM—a closed-world zero-shot classification method—to the open-set OOD setting. Notably, the original GFM model lacks mechanisms specifically designed for OOD detection. To enable OOD detection, we introduce all label names to the GFM and directly use its output logits for zero-shot OOD detection (§3.2). However, despite its effectiveness, the true OOD label space is often inaccessible in real-world settings. Therefore, we investigate zero-shot OOD detection with only ID labels as input (§3.3). This approach, however, suffers from limited expressiveness, as it relies solely on the absence of strong affinity with ID labels to infer OOD status—without any explicit OOD semantic reference, it struggles to detect semantically ambiguous or borderline cases. Therefore, we propose to prompt the LLM to generate a set of pseudo-OOD labels based on the initially unlabeled data, enabling the model to distinguish OOD samples in a more nuanced and fine-grained way. Our method detects OOD nodes by evaluating their affinities with both ID and pseudo-OOD labels (§3.4).
Figure 2: An overview of our framework, GLIP-OOD. Initially, the entire graph is unlabeled, and the labels of all nodes are unknown. GLIP-OOD leverages an LLM to generate semantically informative pseudo-OOD labels from unlabeled data, enabling the GFM to capture nuanced semantic boundaries between ID and OOD classes without requiring any labeled nodes.
3.1 Zero-shot Inference with Graph Foundation Model
Graph Encoder. For each node $v_{i}$, we construct its corresponding induced subgraph $G_{i}=(P_{i},X_{i},A_{i})$, and encode its information as follows:
$g_{i}=g_{\theta}(P_{i},X_{i},A_{i}),$ (1)
where $g_{\theta}$ denotes the Graph Transformer, such as GraphGPS [20], and $P_{i}$ represents the positional embeddings for the subgraph. Applying the mean pooling function yields the graph-level embedding $g_{i}\in\mathbb{R}^{1\times d}$ for the subgraph. In this way, we obtain an embedding $g_{i}$ for each input node $v_{i}$.
Text Encoder. Assuming the graph contains $K$ classes, we can construct a text description for each class label. For instance, in the context of a citation network, a label-associated sentence can be formulated as: “This paper belongs to {class}.” We collect all $K$ such label sentences and denote the set as $\mathcal{T}=\{T_{1},T_{2},\dots,T_{K}\}$. To obtain their embeddings, we apply MiniLM [26] as follows:
$h_{k}=h_{\phi}(T_{k}),$ (2)
where $h_{k}\in\mathbb{R}^{1\times d}$ is the embedding of each label sentence.
Closed-world Zero-shot Inference on Classification Task. The following inference formulation enables zero-shot classification without requiring access to labeled nodes by aligning node embeddings with semantically meaningful label embeddings generated by a text encoder:
$\hat{y}_{i}=\arg\max_{k}\mathbb{E}_{h_{k}}\,\text{sim}(g_{i},h_{k})$ (3)
where $\text{sim}(g_{i},h_{k})$ denotes the similarity between the node embedding $g_{i}$ and the text embedding $h_{k}$ corresponding to the $k$-th label. The label sentence with the highest similarity is selected as the predicted label for node $v_{i}$. However, this approach assumes a closed-world setting, where the set of possible classes is fixed and known in advance. As a result, the model is limited to classifying each node into one of the predefined ID classes. In subsequent sections, we relax this constraint to explore zero-shot open-world scenarios, where OOD nodes may exist beyond the known label set.
3.2 Zero-shot Graph OOD Detection with All Labels
In this section, we provide all the label names to our framework for zero-shot OOD detection. Let $\mathcal{T}_{\text{ID}}=\{T_{1},T_{2},\dots,T_{K_{\text{ID}}}\}$ denote the set of ID label texts, and let $\mathcal{T}_{\text{OOD}}=\{\bar{T}_{1},\bar{T}_{2},\dots,\bar{T}_{K_{\text{OOD}}}\}$ denote the set of OOD label texts. Their corresponding embeddings are denoted as $\mathcal{H}_{\text{ID}}=\{h_{1},h_{2},\dots,h_{K_{\text{ID}}}\}$ and $\mathcal{H}_{\text{OOD}}=\{\bar{h}_{1},\bar{h}_{2},\dots,\bar{h}_{K_{\text{OOD}}}\}$, where $h_{k}=h_{\theta}(T_{k})$ and $\bar{h}_{k}=h_{\theta}(\bar{T}_{k})$ using the same text encoder $h_{\theta}$. We define the OOD score for node $v_{i}$ using its embedding $g_{i}$ and both ID and OOD label embeddings as follows:
$S(v_{i})=S^{*}(\text{sim}(g_{i},\mathcal{H}_{\text{ID}}),\text{sim}(g_{i},\mathcal{H}_{\text{OOD}})),$ (4)
where $\text{sim}(g_{i},\mathcal{H})$ denotes the set of similarity scores between the node embedding $g_{i}$ and the embeddings in the label set $\mathcal{H}$, and $S^{*}$ is a fusion function that computes the OOD score based on these similarities. For example, we can first normalize all the similarity scores (including both ID and OOD labels), and then define the OOD score of node $v_{i}$ as one minus the sum of the normalized similarity scores between the node embedding $g_{i}$ and the sentence embeddings of the ID labels:
$S(v_{i})=1-\sum_{k=1}^{K_{\text{ID}}}\text{sim}(g_{i},h_{k}),$ (5)
where $K_{\text{ID}}$ is the number of ID classes. A higher OOD score indicates lower alignment with any known ID label.
3.3 Zero-shot Graph OOD Detection with ID Labels
Although the previous method is effective, it has a fundamental limitation: it requires the label names of OOD classes. Therefore, we investigate how to perform zero-shot OOD detection using only the ID label names. A simple way to achieve this is to compute the similarity scores between the query node’s embedding $g_{i}$ and only the ID label embeddings, and then normalize these ID similarity scores. These scores can serve as the output logits of an ID classifier, similar to those used in supervised graph learning methods. Therefore, we can apply any post-hoc OOD detector—such as MSP [9] or the Energy score [14]—to these ID similarity scores to perform OOD detection.
Another simple method is to use a set of negative prompts to substitute the OOD label sentences in order to capture the negation semantics with nodes [24]. Specifically, instead of comparing a node’s embedding with known OOD labels, we measure its similarity to a set of templated negative prompts such as “This node does not belong to [ID class]” or “This is not about [ID label].” These prompts are designed to express semantic exclusion and allow the model to implicitly learn what it means for a node to fall outside the ID classes. By contrasting a node’s affinity to ID label prompts and these negative prompts, we estimate its likelihood of being OOD without requiring access to the actual OOD label space. In this case, these negative prompts form the set of pseudo-OOD label texts $\mathcal{T}_{\text{OOD}}=\{\bar{T}_{1},\bar{T}_{2},\dots,\bar{T}_{K_{\text{OOD}}}\}$. Given $\mathcal{T}_{\text{ID}}$ and $\mathcal{T}_{\text{OOD}}$, we can use Eqn. 4 to compute OOD scores for any input node. However, we find that these two simple methods do not perform well in practice. Therefore, we need to design more effective approaches to enhance the GFM’s OOD awareness.
3.4 Zero-shot Graph OOD Detection with Pseudo-OOD Labels
OOD label names generation with LLM. To enhance the model’s OOD awareness, we aim to obtain negative labels that are semantically distant from the given ID labels but close to the actual, unknown OOD labels. Since all existing node-level graph OOD detection methods operate in a transductive setting, we can access the full set of nodes during training, but not their labels. Therefore, as shown in Fig. 2, we propose using LLMs to generate pseudo-OOD label names based on the initially unlabeled graph by leveraging the textual information of the nodes. We first randomly sample a small set of nodes and prompt the LLM to perform zero-shot open-world annotation. Specifically, we provide the names of the ID classes to the LLM and ask whether each unlabeled query node belongs to any of these classes, based on its textual information. The LLM is prompted to determine:
$\hat{y}_{i}\;=\;\operatorname{LLM}\!\bigl{(}v_{i},\mathcal{Y}_{\text{ID}}\bigr{)}\;:=\;\begin{cases}y\in\mathcal{Y}_{\text{ID}},&\text{if the LLM matches }v_{i}\text{ to an ID class},\\[4.0pt] \text{OOD},&\text{otherwise}.\end{cases}$ (6)
If the LLM predicts that a node $v_{i}$ does not belong to any ID category, we treat it as a pseudo-OOD node. We then prompt the LLM to summarize the node’s textual content and generate a suitable OOD category name for it.
[…]
OOD detection in augmented label space. Once a set of pseudo-OOD label names is generated by the LLM, we construct an augmented label space that includes both the original ID classes and the newly synthesized OOD labels. The generated OOD categories form the set $\mathcal{Y}_{\text{OOD}}$, and the augmented label space is defined as:
$\mathcal{Y}_{\text{aug}}=\mathcal{Y}_{\text{ID}}\cup\mathcal{Y}_{\text{OOD}}.$ (8)
We encode each label in $\mathcal{Y}_{\text{aug}}$ into a text embedding using the pretrained text encoder, as shown in Eqn. 2. For each node, we compute the similarity between its embedding and the embeddings of all labels in the augmented label space. A node is classified as OOD if it has low affinity to all ID label embeddings but exhibits high similarity to one or more pseudo-OOD labels. This approach allows the model to leverage both the discriminative power of ID labels and the semantic coverage of LLM-generated OOD labels, enabling more accurate and fine-grained OOD detection without requiring access to any real OOD label names.
4 Experiments
4.1 Experimental Setup
Datasets We utilize the following TAG datasets, which are commonly used for node classification: Cora [16], Citeseer [7], Ele-Computers [30] and Wiki-CS [17]. The reason for using these datasets is that we perform zero-shot OOD detection based on GraphCLIP [33], and these datasets were not used during the pretraining of this GFM. Therefore, we can effectively evaluate its zero-shot OOD detection performance on unseen data. For each dataset, we split all classes into ID and OOD sets and the ID classes for the four datasets are shown in Appendix B. Additionally, the number of ID classes is set to a minimum of three to perform ID classification task.
Baselines We evaluate GLIP-OOD against two categories of baselines, including: (1) Supervised graph OOD detection methods, including MSP [9], Entropy, Energy [14], GNNSafe [27], and GRASP [15]. (2) Language model methods for zero-shot OOD detection, including BERT [4], SBERT [21], DeBERTa [8], and E5 [25].
Settings For all methods that require labeled nodes to train the ID classifier, we follow the setting in [22] and use $20\times K$ labeled ID nodes for training on each dataset with $K$ ID classes. Additionally, we randomly select $10\times K$ labeled ID nodes and an equal number of OOD nodes for validation. The test set consists of 500 randomly selected ID nodes and 500 OOD nodes. For language model-based methods, we use them to encode the text associated with nodes and the information of ID labels, respectively. We then compute similarity scores between each node embedding and the embeddings of ID label sentences. Based on the normalized similarity scores, we calculate the MSP score, which is used as the OOD score. For GLIP-OOD, we use GPT-4o-mini to identify pseudo-OOD nodes and generate corresponding pseudo-OOD label names. The prompts used in this process are provided in Appendix D. Our method has two variants: (1) using all real labels, and (2) using both ID labels and pseudo-OOD labels to perform zero-shot OOD detection.
Implementation details For all supervised methods, we employ a two-layer GCN as the ID classifier and apply various post-hoc OOD detectors to it. Each GCN is configured with a hidden dimension of 32, a learning rate of 0.01, a dropout rate of 0.5 and a weight decay of 5e-4. For all experiments, results are averaged over 5 different random seeds. All experiments are conducted on hardware equipped with an NVIDIA RTX 4080 Ti.
Evaluation Metrics For the ID classification task, we use classification accuracy (ACC) as the evaluation metric. For the OOD detection task, we employ three commonly used metrics from the OOD detection literature [27]: the area under the ROC curve (AUROC), the precision-recall curve (AUPR), and the false positive rate when the true positive rate reaches 95% (FPR@95). In all experiments, OOD nodes are considered positive cases.
4.2 Main Results
The OOD detection results are shown in Table 1. From the results, we conclude the following: (1) Exposing all class labels to GLIP-OOD enables it to achieve the best performance among all methods, even surpassing fully supervised approaches that rely on a large number of ID labels to train the ID classifier. This demonstrates that the GFM is inherently strong enough to perform zero-shot OOD detection and has significant potential to become highly OOD-aware, provided the OOD classes are clearly defined. (2) For language model baselines performing zero-shot OOD detection, relying solely on ID class labels is inadequate for distinguishing ID and OOD nodes in most cases. (3) Using our proposed approach to generate pseudo-OOD labels significantly improves the OOD awareness of the GFM, compared to using only ID class labels to compute similarity scores and applying post-hoc OOD detector on top of the language model’s output logits. (4) When using ID labels and pseudo-OOD labels, GLIP-OOD achieves performance comparable to that of supervised methods. However, GLIP-OOD requires neither labeled nodes for training nor any knowledge of OOD labels. The only information it needs is the names of the ID classes, after which it can directly perform zero-shot OOD detection for any input node. (5) While GLIP-OOD is primarily designed for OOD detection, it also performs competitively in ID classification. Although it lags slightly behind fully supervised GNN-based classifiers in terms of accuracy (as expected given the absence of node-level supervision), its performance remains strong enough to support reliable OOD detection.
Table 1: Performance comparison (best OOD detection results highlighted in bold) of different models on ID classification and OOD detection tasks.
Method Models Cora Citeseer Ele-Computers Wiki-CS
ACC $\uparrow$ AUROC $\uparrow$ AUPR $\uparrow$ FPR95 $\downarrow$ ACC $\uparrow$ AUROC $\uparrow$ AUPR $\uparrow$ FPR95 $\downarrow$ ACC $\uparrow$ AUROC $\uparrow$ AUPR $\uparrow$ FPR95 $\downarrow$ ACC $\uparrow$ AUROC $\uparrow$ AUPR $\uparrow$ FPR95 $\downarrow$
Supervised GNN MSP 0.9284 0.7831 0.7893 0.7832 0.8620 0.7316 0.7232 0.7868 0.8452 0.6518 0.6644 0.9356 0.8744 0.7884 0.7849 0.7020
Entropy 0.9280 0.7853 0.7883 0.7740 0.8620 0.7425 0.7289 0.7392 0.8416 0.6583 0.6660 0.9236 0.8704 0.7987 0.7931 0.6720
Energy 0.9280 0.8002 0.8025 0.7664 0.8616 0.7727 0.7605 0.7016 0.8496 0.6329 0.6518 0.9352 0.8720 0.8126 0.8095 0.6742
GNNSafe 0.9272 0.8718 0.8802 0.5860 0.8604 0.7836 0.7581 0.7604 0.8456 0.7438 0.7689 0.8312 0.8732 0.8925 0.9039 0.5044
GRASP 0.9248 0.8460 0.8379 0.6020 0.8596 0.7457 0.7178 0.7844 0.8456 0.6868 0.7247 0.8608 0.8760 0.8958 0.8979 0.4532
Language Models BERT 0.5328 0.5028 0.4877 0.9296 0.5992 0.5759 0.5422 0.9092 0.4916 0.4064 0.4558 0.9908 0.5236 0.5234 0.5022 0.8808
SBERT 0.8436 0.7381 0.6730 1.0000 0.8044 0.6823 0.6441 1.0000 0.5728 0.6045 0.5640 0.7824 0.7312 0.6622 0.6199 0.8116
DeBERTa 0.5920 0.5282 0.5113 0.9184 0.4512 0.4294 0.4452 0.9272 0.3068 0.4869 0.5004 0.9536 0.5564 0.6177 0.6000 0.8928
E5 0.8112 0.7868 0.7390 0.5620 0.7700 0.7035 0.6791 0.7964 0.5952 0.5293 0.5086 0.8588 0.7004 0.7074 0.6673 0.7596
Ours GLIP-OOD-L 0.8836 0.7573 0.7199 0.7228 0.7708 0.7963 0.7892 0.9248 0.7932 0.7265 0.6897 0.6032 0.8244 0.8204 0.8137 0.6420
GLIP-OOD-R 0.8836 0.8810 0.8542 0.4176 0.7708 0.8745 0.8820 0.5584 0.7932 0.8742 0.8646 0.4836 0.8244 0.9282 0.9257 0.3476
Note: GLIP-OOD-L denotes the use of labels generated by the LLM, GLIP-OOD-R denotes the use of real labels.
4.3 Visualization of Label Space
We plot the embedding space of class label representations on the Ele-Computers and Wiki-CS datasets, where ID labels, real OOD labels, and LLM-generated pseudo-OOD labels are visualized within the same semantic space. The pseudo-OOD labels, generated using the method described in Section 3.4, are further clustered based on their semantic similarity. We instruct the LLM to directly output the names of the clustered labels. The corresponding prompt is detailed in Appendix D. From the results in Fig. 3, we observe that the pseudo-OOD labels generated by LLMs lie outside the ID cluster but remain relatively close to real OOD labels, indicating that they serve as effective proxies. This provides semantic evidence that LLM-generated pseudo-OOD labels occupy intermediate and informative regions in the embedding space. These labels help expand the model’s decision boundaries in a controlled manner, thereby improving zero-shot OOD detection performance.
Refer to captionRefer to caption
Figure 3: Visualization of the semantic label space for the Ele-Computers and Wiki-CS dataset. Each point represents a label name embedded in the GLIP-OOD’s semantic space.
4.4 More OOD Scores
We use additional OOD scoring functions for zero-shot OOD detection with all labels. Following Eqn. 5, the other three OOD scores are defined as follows:
[…]
Table 2: Comparison of OOD detection performance across different OOD scoring methods.
Methods Cora Citeseer Ele-Computers Wiki-CS
AUROC $\uparrow$ AUPR $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$ AUPR $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$ AUPR $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$ AUPR $\uparrow$ FPR95 $\downarrow$
sum_gap 0.8826 0.8593 0.4176 0.8756 0.8833 0.5592 0.8742 0.8645 0.4836 0.9281 0.9255 0.3476
max_gap 0.8713 0.8511 0.4284 0.8706 0.8802 0.5840 0.8690 0.8595 0.5312 0.9150 0.9166 0.4284
OOD_ratio 0.8879 0.8847 0.4180 0.8776 0.8924 0.5556 0.8743 0.8677 0.4836 0.9286 0.9280 0.3476
sum_ID 0.8810 0.8542 0.4176 0.8745 0.8820 0.5584 0.8742 0.8646 0.4836 0.9282 0.9257 0.3476
From the results in Table 2, we observe that all four post-hoc OOD scores generally yield similar OOD detection performance, and GLIP-OOD remains robust across different OOD scoring functions.
4.5 Zero-shot OOD Detection with Weak OOD Information
In this experiment, we evaluate the performance of GLIP-OOD under a zero-shot setting where the model is provided with all class label names. This represents a weak OOD information scenario, where OOD label names are known but no node-level supervision is available. We compare our method against a variety of language model baselines, including BERT, SBERT, DeBERTa, and E5, which compute similarity scores between each node’s text and the label names for OOD detection.
As shown in Table 3, GLIP-OOD consistently outperforms all language models across all datasets. This highlights the effectiveness of the GFM; even without training on node labels, the GFM can effectively leverage label names and graph structure to learn semantically meaningful representations, enabling robust OOD detection in a fully zero-shot setting.
Table 3: Performance comparison (best results highlighted in bold) of different models on zero-shot OOD detection when provided with all label names.
Models Cora Citeseer Ele-Computers Wiki-CS
AUROC $\uparrow$ AUPR $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$ AUPR $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$ AUPR $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$ AUPR $\uparrow$ FPR95 $\downarrow$
BERT 0.6895 0.6534 0.7772 0.7378 0.7266 0.7372 0.5266 0.5023 0.8732 0.7332 0.7577 0.8268
SBERT 0.8511 0.8168 0.4748 0.8662 0.8763 0.5724 0.7712 0.7490 0.6628 0.8810 0.8809 0.5056
DeBERTa 0.6785 0.6528 0.7948 0.6743 0.6431 0.7884 0.6920 0.6327 0.7172 0.6694 0.6918 0.9000
E5 0.8636 0.8474 0.4700 0.8295 0.8302 0.6372 0.6869 0.6536 0.7312 0.8539 0.8555 0.5672
GLIP-OOD 0.8810 0.8542 0.4176 0.8745 0.8820 0.5584 0.8742 0.8646 0.4836 0.9282 0.9257 0.3476
4.6 More OOD and ID Class Divisions
Table 4: Comparison of OOD detection performance (best results highlighted in bold) across different models under various OOD and ID class divisions.
Methods Cora Citeseer Ele-Computers Wiki-CS
AUROC $\uparrow$ AUPR $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$ AUPR $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$ AUPR $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$ AUPR $\uparrow$ FPR95 $\downarrow$
MSP 0.7982 0.7786 0.7436 0.7583 0.7667 0.8208 0.7305 0.7370 0.8320 0.7737 0.7739 0.7796
GNNSafe 0.8819 0.8831 0.5992 0.8029 0.8142 0.8300 0.8342 0.8234 0.5432 0.8646 0.8755 0.6292
GRASP 0.8891 0.8886 0.5844 0.7612 0.7540 0.8408 0.8087 0.7758 0.5656 0.8696 0.8500 0.5116
GLIP-OOD 0.9273 0.9185 0.3484 0.9253 0.9182 0.7016 0.8919 0.8975 0.4952 0.9328 0.9299 0.3008
To further evaluate the robustness and generalizability of GLIP-OOD, we experiment with an alternative set of OOD and ID class divisions across four datasets. The new ID classes are listed in Appendix B. From the results in Table 4, we observe that GLIP-OOD consistently achieves the best performance across all datasets and metrics. This demonstrates that the GFM possesses strong inherent capability for OOD detection and can adapt to different open-set configurations, thereby enhancing both the reliability and robustness of zero-shot graph OOD detection.
5 Related Works
5.1 Graph OOD Detection
Node-level graph OOD detection aims at detecting nodes whose attributes or structural context deviate from the distribution of known ID classes, despite being embedded in the same transductive graph during training and inference. GNNSafe [27] demonstrates that standard GNN classifiers inherently exhibit OOD detection capabilities and introduces an energy-based discriminator trained using conventional classification loss. OODGAT [22] explicitly distinguishes inliers from outliers during feature propagation and solves the problem of node classification and outlier detection in a joint framework. GRASP [15] explores the advantages of propagating OOD scores and provides theoretical insights into the conditions under which this propagation is effective. Furthermore, it proposes an edge augmentation strategy with theoretical guarantees to enhance post-hoc node-level OOD detection. However, all existing graph OOD detection methods rely on a large number of labeled ID nodes for training. Our method is the first to require no labeled nodes and is the first approach for zero-shot graph OOD detection.
5.2 Zero-shot OOD Detection
Zero-shot OOD detection has been widely studied in the image domain [24, 5, 18, 1] and text domain [32, 10]. These methods typically provide large-scale pretrained models with only the label names of ID classes and use similarity scores to distinguish ID from OOD inputs. For example, EOE [1] designs LLM prompts guided by visual similarity to generate potential outlier class labels tailored for zero-shot OOD detection. ZOC [5] extends the CLIP model to dynamically generate candidate unseen labels for each test image and introduces a novel confidence score based on the similarity between the test image and both seen and generated unseen labels in the feature space. CLIPN [24] enhances CLIP with the ability to distinguish between ID and OOD samples by leveraging positive-semantic and negation-semantic prompts. [10] proposes an ID-label-free OOD detection method that leverages unsupervised clustering and contrastive learning to learn effective data representations for OOD detection. Despite these advances, zero-shot OOD detection in graph domains remains largely unexplored. This is due to the unique challenges posed by relational inductive biases and the lack of large-scale, general-purpose foundation models specifically trained for graphs. Our work bridges this gap by enabling zero-shot node-level OOD detection in graphs through the use of GFMs and language models for pseudo-OOD label generation—without requiring any labeled nodes.
6 Conclusion
In this work, we introduce GLIP-OOD, a novel framework for zero-shot graph OOD detection that leverages the capabilities of graph foundation model. Unlike existing graph OOD detection methods that rely on large numbers of labeled ID nodes, GLIP-OOD operates in a fully zero-shot setting, requiring only the names of ID classes and the graph structure. By prompting LLMs to generate pseudo-OOD labels, our method enables GFMs to learn meaningful semantic distinctions between ID and OOD nodes without any supervised training. Through extensive evaluation on four benchmark datasets, we demonstrate that when provided with all label names, GLIP-OOD outperforms all supervised methods. When given only ID label names, it achieves performance comparable to supervised approaches that rely on a large number of labeled ID nodes for training. Future research could explore more effective strategies for generating negative labels to further enhance the inherent OOD-awareness of increasingly powerful graph foundation models.
Broader Impact By enhancing the OOD-awareness of graph-based systems, our method advances the development of safer and more trustworthy AI models for open-world environments. It also paves the way for future research at the intersection of foundation models, LLMs, and graph OOD detection, offering a promising step toward scalable, reliable, and label-efficient open-world graph learning.
Limitations While our proposed method eliminates the need for both labeled nodes and a training process for graph OOD detection, it is applicable only to TAGs. Additionally, reliance on LLMs for label generation introduces concerns regarding the biases present in the pretraining of those models.
Appendix A Dataset
Cora
The Cora dataset [16] consists of 2,708 scientific publications, each classified into one of seven research topics: case-based reasoning, genetic algorithms, neural networks, probabilistic methods, reinforcement learning, rule learning, and theory. Nodes represent individual papers, while edges denote citation relationships, resulting in a graph structure with 5,429 edges.
CiteSeer
The CiteSeer dataset [7] comprises 3,186 scientific articles categorized into six research domains: Agents, Machine Learning, Information Retrieval, Databases, Human-Computer Interaction, and Artificial Intelligence. Each node corresponds to a paper, with features derived from its title and abstract. Citation relationships between papers define the edges of the graph.
WikiCS
Wiki-CS [17] is designed for benchmarking graph neural networks. Each node represents a computer science article, categorized into one of ten subfields used as classification labels. Edges correspond to hyperlinks between articles, and node features are extracted from the text content of the articles.
Ele-Computer
Ele-Computer dataset [30] consists of nodes representing electronic products, with edges capturing frequent co-purchases or co-views. Each node is labeled based on a hierarchical three-level classification of electronics. The associated text attribute is either the most upvoted user review or a randomly selected one if no such review exists. The classification task aims to assign each product to one of ten predefined categories.
Appendix B ID Classes
Table 5: ID classes for different datasets.
Dataset ID class
Cora [0, 1, 2]
Citeseer [0, 1, 2]
Wiki-CS [0, 1, 2, 3]
Ele-Computers [0, 1, 2]
Table 6: ID classes for each dataset in another split.
Dataset ID class
Cora [2, 3, 4]
Citeseer [2, 3, 4]
Wiki-CS [2, 3, 4, 5]
Ele-Computers [2, 3, 4, 5]
Appendix C More Results
Table 7: Comparison of OOD detection performance on Cora and Citeseer across different language models using ID labels.
Models Post-hoc Cora Citeseer
AUROC AUPR FPR95 AUROC AUPR FPR95
BERT Energy 0.5377 0.5104 0.9000 0.6340 0.6069 0.8844
Entropy 0.4963 0.4820 0.9284 0.5676 0.5314 0.9092
SBERT Energy 0.7605 0.6966 0.6420 0.7754 0.7317 0.6952
Entropy 0.7603 0.6891 0.5920 0.6912 0.6601 0.7632
DeBERTa Energy 0.5324 0.5107 0.9324 0.5642 0.5518 0.9080
Entropy 0.5047 0.5028 0.9244 0.4465 0.4504 0.9268
E5 Energy 0.7503 0.6895 0.6580 0.7540 0.7230 0.7636
Entropy 0.7931 0.7477 0.5584 0.7201 0.7013 0.7944
Table 8: Comparison of OOD detection performance on Ele-Computers and Wiki-CS across different language models using ID labels.
Models Post-hoc Ele-Computers Wiki-CS
AUROC AUPR FPR95 AUROC AUPR FPR95
BERT Energy 0.4356 0.4717 0.9692 0.4934 0.4964 0.9428
Entropy 0.4277 0.4682 0.9908 0.5176 0.4935 0.8780
SBERT Energy 0.5797 0.5431 0.8556 0.6668 0.5932 0.7228
Entropy 0.6031 0.5600 0.7816 0.6669 0.6332 0.7984
DeBERTa Energy 0.4804 0.4895 0.9460 0.4777 0.4880 0.9572
Entropy 0.6538 0.6190 0.8808 0.6423 0.6218 0.8844
E5 Energy 0.5513 0.5145 0.8692 0.5445 0.5097 0.8724
Entropy 0.5192 0.4997 0.8568 0.7240 0.6854 0.7596
Appendix D Prompts
In our experiments, we first randomly sample 80 nodes from each of the Cora and Citeseer datasets, and 120 nodes from each of the Ele-Computers and Wiki-CS datasets. For each sampled node, we replace the placeholder “Object Content” in the prompt templates with its associated text, and then invoke the LLM to classify it into one of the predefined ID classes or, if necessary, generate a new class name. Optionally, we can filter out low-quality pseudo-OOD labels generated by the LLM. During zero-shot inference, we randomly select a small subset of the remaining labels to form the pseudo-OOD label set and perform OOD detection using both these pseudo-OOD labels and the original ID labels.
Zero-Shot Prompt for Class Name Generation for Ele-Computers You are a highly knowledgeable text-classification system. You have these in-distribution (ID) classes: Category 1 Category 2 Dataset description: Ele-Computers dataset is extracted from the Amazon Electronics dataset. Below is the text of a Paper: [Insert Object Content Here] Your task is to classify this Paper based on the following instructions: 1. If this Paper does not belong to any of the listed ID classes: Use your general knowledge to determine the single most appropriate class name of an electronic product that is not already among the ID classes listed above. The new class should not be too specific nor too detailed, and it should be an electronic product commonly seen. You may internally summarize or reason through the text. Your final output must be only the name of the generated class. 2. If this Paper does belong to one of the listed ID classes: Your final output must be the exact name of that ID class listed above. Important Notes: Do not select any class from the ID list unless the Paper clearly fits. Your response must be a single line containing only the class name — no punctuation, no extra text. Answer:
Zero-Shot Prompt for Class Name Generation for Cora You are a highly knowledgeable text-classification system. You have these in-distribution (ID) classes: Category 1 Category 2 Dataset description: Cora dataset consists of scientific publications. Below is the text of a Paper: [Insert Object Content Here] Your task is to classify this Paper based on the following instructions: 1. If this Paper does not belong to any of the listed ID classes: Use your general knowledge to determine the single most appropriate class name of a machine learning-related topic that is not already among the ID classes listed above. The new class should not be too specific nor too detailed, and it should be a commonly-seen broad topic in the machine learning-related field — avoid rare, niche, or overly obscure concepts. You may internally summarize or reason through the text. Your final output must be only the name of the generated class. 2. If this Paper does belong to one of the listed ID classes: Your final output must be the exact name of that ID class listed above. Important Notes: Do not select any class from the ID list unless the Paper clearly fits. Your response must be a single line containing only the class name — no punctuation, no extra text. Answer:
General Zero-Shot Prompt for Class Name Generation for Citeseer You are a highly knowledgeable text-classification system. You have these in-distribution (ID) classes: Category 1 Category 2 Below is the text of a Paper: [Insert Object Content Here] Your task is to classify this Paper based on the following instructions: 1. If this Paper does not belong to any of the listed ID classes: Use your general knowledge to determine the single most appropriate class name that is not already among the ID classes listed above. You may internally summarize or reason through the text. Your final output must be only the name of the new class. 2. If this Paper does belong to one of the listed ID classes: Your final output must be the exact name of that ID class listed above. Important Notes: Do not select any class from the ID list unless the Paper clearly fits. Your response must be a single line containing only the class name — no punctuation, no extra text. Answer:
Zero-Shot Prompt for Class Name Generation for Wiki-CS You are a highly knowledgeable text-classification system. You have these in-distribution (ID) classes: Category 1 Category 2 Dataset description: WikiCS dataset is a Wikipedia-based dataset, comprising many computer science branches as classes characterized by high connectivity. Node features are extracted from the corresponding article texts. Below is the text of a Paper: [Insert Object Content Here] Your task is to classify this Paper based on the following instructions: 1. If this Paper does not belong to any of the listed ID classes: Use your general knowledge to determine the single most appropriate class name of a topic in computer science and engineering that is not already among the ID classes listed above. The new class should not be too specific nor too detailed, and it should be a commonly-seen broad topic — avoid rare, niche, or overly obscure concepts. You may internally summarize or reason through the text. Your final output must be only the name of the generated class. 2. If this Paper does belong to one of the listed ID classes: Your final output must be the exact name of that ID class listed above. Important Notes: Do not select any class from the ID list unless the Paper clearly fits. Your response must be a single line containing only the class name — no punctuation, no extra text. Answer:
For the Ele-Computers dataset, we cluster the generated labels into higher-level categories using the procedure described in the following prompt. These clustered categories are then used as pseudo-OOD labels in the subsequent steps.
Clustering Prompt for Category Grouping You are a highly intelligent taxonomy generation system. You are given a list of items to be clustered: generated class 1 generated class 2 Task: cluster the above things into {num_cluster} categories and specify their names, requests: The generated categories should be away from the below categories: in-distribution class 1 in-distribution class 2 please generate one or two sentences to describe the above {num_cluster} things, respectively. Answer: